### v1

# Self-Rewarding Vision-Language Model via Reasoning Decomposition

Abstract
Vision-Language Models (VLMs) often suffer from visual hallucinations – saying things that aren’t actually in the image – and language shortcuts, where they skip the visual part and just rely on text priors. These issues arise because most post-training methods for VLMs rely on simple verifiable answer matching and supervise only final outputs, leaving intermediate visual reasoning without explicit guidance. As a result, VLMs receive sparse visual signals and often learn to prioritize language-based reasoning over visual perception. To mitigate this, some existing methods add visual supervision using human annotations or distilled labels from external large models. However, human annotations are labor-intensive and costly, and because external signals cannot adapt to the evolving policy, they cause distributional shifts that can lead to reward hacking.
In this paper, we introduce Vision-SR1, a self-rewarding method that improves visual reasoning without relying on external visual supervisions via reinforcement learning. Vision-SR1 decomposes VLM reasoning into two stages: visual perception and language reasoning. The model is first prompted to produce self-contained visual perceptions that are sufficient to answer the question without referring back the input image. To validate this self-containment, the same VLM model is then re-prompted to perform language reasoning using only the generated perception as input to compute reward. This self-reward is combined with supervision on final outputs, providing a balanced training signal that strengthens both visual perception and language reasoning. Our experiments demonstrate that Vision-SR1 improves visual reasoning, mitigates visual hallucinations, and reduces reliance on language shortcuts across diverse vision-language tasks.
Figure: (Left) Case study comparing Vision-R1 with our method, showing reduced hallucinations. (Right) Reinforcement learning training curve compared to Vision-R1, along with accuracy on three widely used vision-language benchmarks.
[Uncaptioned image] Code: https://github.com/zli12321/Vision-SR1.
1 Introduction
Recent advances in vision-language models (VLMs) have progressed by integrating pre-trained language models and vision encoders with instruction tuning (Liu et al., 2023b; et al, 2024; Chen et al., 2024; Bai et al., 2025; Li et al., 2025d). Despite these successes, a critical limitation remains in their reasoning capabilities: VLMs often produce visual hallucinations – descriptions of content that is not actually present in the image (Guan et al., 2024; Liu et al., 2024; Li et al., 2025e; Liu et al., 2023a) – or rely on language shortcuts, where the model bypasses visual understanding and instead depends solely on text priors (Si et al., 2022; Bleeker et al., 2024). Very recently, R1-style reinforcement learning (RL) methods have been shown to improve the reasoning abilities of VLMs across diverse tasks (Huang et al., 2025; Shen et al., 2025; Xia et al., 2025; Zhang et al., 2025). However, these methods often encourage “thinking over seeing” that lean heavily on language reasoning while underutilize visual perception (Liu et al., 2025; Yao et al., 2025). This imbalance makes VLMs susceptible to reward hacking (Fu et al., 2025) and spurious effects (Shao et al., 2025) observed in RL training. Although VLMs trained with RL often have apparent improvements, they can largely reflect probability shifts toward the style of training and test data, leading to language shortcut answers from prior knowledge and overlooking hallucination risks (Li et al., 2025b).
In essence, most existing post-training methods for VLMs rely on simple verifiable answer matching and thus lack explicit supervision of visual information. As a result, VLMs receive sparse visual signals and often learn to prioritize language-based reasoning over visual perception. To mitigate this, some methods introduce intermediate visual supervision through human annotations (Thawakar et al., 2025) or distilled labels (e.g., pre-extracted key steps) from external models (Xu et al., 2024; Zhang et al., 2025; Xiao et al., 2025; Xia et al., 2025; Lu et al., 2025). However, these solutions face significant limitations. Human annotations are labor-intensive, costly, and difficult to scale across multimodal tasks, while distilled signals inherit biases from source models and often fail to generalize across diverse domains. Moreover, distributional shifts between fixed intermediate signals and the continually updated policy can lead to reward hacking (Gao et al., 2023). Most importantly, both approaches still rely on external supervision, restricting their scalability and applicability.
In this paper, we introduce Vision-SR1, a reinforcement learning framework that encourages VLM to produce self-contained visual reasoning that can be verified by the VLM itself, without external supervision. Vision-SR1 decomposes the reasoning process into two stages: visual perception and language reasoning. The visual perception is required to capture all details relevant to answering the query, so that the reasoning stage can proceed without re-accessing the original image.
The training involves two rollout passes of the same VLM:
[…]
  • The model generates a structured output that explicitly separates visual perception, chain-of-thought (CoT) reasoning, and the final answer.
  • An accuracy reward is computed by comparing the final answer with the ground truth.
[…]
  • The model is re-prompted to reason using only the generated perception (without re-accessing the original image). If the correct answer is derived, the perception is considered faithful, and a self-visual reward is assigned.
Both rewards are combined to provide a balanced training signal that strengthens both visual perception and language reasoning. Experiments show that Vision-SR1 improves visual reasoning, mitigates hallucinations, and reduces language shortcuts across diverse vision-language tasks.
2 Method
We build upon advancement of multimodal Group Relative Policy Optimization (GRPO) (Shao et al., 2024) for improving VLM reasoning. We first review the key concepts then introduce our method.
2.1 Preliminary: Reinforcement Learning for VLM with Verifiable Reward
We denote a pre-trained VLM as a policy model $\pi$ to be optimized in reinforcement learning. Given a multimodal question ($Q$) consists of an image $i$ and a text question $q$, where $Q=\{i,q\}$, the policy model $\pi$ generates a reasoning response $s$. We use GRPO to optimize the response $s$ for the policy model. For each multimodal question $Q=\{i,q\}$ we sample a group of $K$ candidate responses $\mathcal{S}_{Q}=\{s_{1},\dots,s_{K}\},\;s_{k}\sim\pi_{\theta}(\,\cdot\,|Q).$ Each response is scored by a scalar reward $r(Q,s_{k})$ (defined in Sec. 2.2), and we compute a group‑relative advantage
$\hat{A}^{\text{grp}}(Q,s_{k})\;=\;r(Q,s_{k})\;-\;\frac{1}{K}\sum_{j=1}^{K}r(Q,s_{j}),$ (1)
which centres rewards within the group, removing question‑level biases while retaining pairwise preferences. We update the policy by maximizing
$\mathcal{L}_{\text{GRPO}}(\theta)\;=\;\mathbb{E}_{Q\sim\mathcal{D}}\Bigl{[}\sum_{k=1}^{K}\hat{A}^{\text{grp}}(Q,s_{k})\;\log\pi_{\theta}(s_{k}\,|\,Q)\;-\;\beta\,\operatorname{KL}\!\bigl{(}\pi_{\theta}(\,\cdot\,|Q)\;\|\;\pi_{\theta_{0}}(\,\cdot\,|Q)\bigr{)}\Bigr{]},$ (2)
where $\pi_{\theta_{0}}$ is the frozen, pre‑trained reference model and $\beta$ controls the strength of the KL penalty that keeps the updated policy close to its original behavior.
The group‑centred baseline in (1) guarantees $\sum_{k}\hat{A}^{\text{grp}}(Q,s_{k})=0$, thereby reducing the variance of policy‑gradient estimates without requiring an external value critic.
2.2 Our Method: Self-Rewarding VLM via Reasoning Decomposition
As we discussed, incorporating intermediate visual supervision can strengthen the reasoning ability of VLMs. However, existing methods suffer from key limitations: methods based on human annotations are labor-intensive and costly (Thawakar et al., 2025), while approaches that distill supervision from external models provide static signals that cannot adapt as the policy model itself evolves during training (Zhang et al., 2025; Xiao et al., 2025; Xia et al., 2025). To overcome these issues, we introduce a self-rewarding framework that enables the VLM to reward its own visual perception. The key idea is to decompose the visual reasoning process into structured components, i.e., the VLM first produces a self-contained visual perception and then assesses whether this perception is sufficient for produce the final answer. This decomposition reduces reliance on external supervision and allows the reward signal to adapt dynamically as the model improves.
Decomposed VLM Reasoning.
To encourage the VLM to perform self-contained visual reasoning, we require every response to adhere to a See-Think generation format (Jia et al., 2024; Xia et al., 2025) format. Specifically, for a vision-language task, $Q=\{i,q\}$ where $i$ is the input image and $q$ is the textual query, the model produces the following structured output:
$\text{{<visual perception{}>}}\,c\;\text{{</visual perception{}>}}\;\|\;\text{{<think>}}\,t\;\text{{</think>}}\;\|\;\text{{<answer>}}\,a\;\text{{</answer>}},$
where $c$ is a self-contained visual perception that captures all visual information necessary to solve the task, so that the following language reasoning can proceed without re-accessing the original input image. Besides, $t$ is the language reasoning trace, and $a$ denotes the final answer.
Self-Reward for Visual Perception.
A challenge is judging whether the visual perception $c$ is self-contained – i.e. whether it encodes all the visual information needed to answer the question $Q=\{i,q\}$ correctly. To address this, our idea is to treat the visual perception as a text-only proxy for the image and validate it by re-prompting the VLM itself to perform language reasoning using only the generated perception as input. If the model can derive the correct answer from $(c,q)$ alone, we consider $c$ to be visually faithful and assign the corresponding visual reward.
$\hat{a}\;=\;f_{\theta}\bigl{(}c,q\bigr{)},\quad r_{\mathrm{visual}}(Q,c)\;=\;\mathbb{I}\!\bigl{[}\hat{a}=a^{*}\bigr{]},$ (3)
where $a^{*}$ is the ground-truth answer. Instead of using an external reward model, we leverage the policy model’s own reasoning ability for self-evaluation. The model itself determines the reward by answering the question using only its generated visual perception (Figure 1).
Total Reward.
The total reward comprises three aligned components, each conditioned on the question $Q=\{i,q\}$:
  • Format reward $r_{\mathrm{fmt}}(s)$: measures whether the response strictly follows the required layout.
  • Accuracy reward $r_{\mathrm{ans}}(Q,a)$: measures the correctness of the final answer. Because $a$ is generated after the reasoning trace $t$, the term implicitly rewards CoT reasoning.
  • Visual reward $r_{\mathrm{cap}}(Q,c)$: measures whether the visual perception output is self-contained, i.e., sufficient to answer the question without image. A reward of 1 is assigned if, given only the question and the visual perception, the VLM can give the correct answer.
The entangling scalar reward is weighted as
$r(Q,s)=\,r_{\mathrm{visual}}(Q,c)+\,r_{\mathrm{ans}}(Q,a),+\alpha\,r_{\mathrm{fmt}}(s)$ (4)
with hyper-parameters $(0\leq\alpha\leq 1)$ tuned empirically. The answer accuracy reward $r_{\mathrm{ans}}(Q,a)$ is computed on the final answer $a$, produced autoregressively after the visual perception $c$. Because the decoder conditions on $c$, this term implicitly encourages coherent reasoning and tight coupling between visual perception and textual reasoning.
Figure 1: Overall framework of Vision-SR1. During RL training, the VLM performs two rollouts. In the first pass, the model takes an image–query pair and generates a structured output (visual perception, CoT reasoning, and answer), with answer reward computed against the ground truth. In the second pass, the model is re-prompted to answer using only query and its generated visual perception. If the correct answer is derived, a self-visual reward is assigned.
2.3 Theoretical Analysis
We analyze why incorporating a perception-level reward improves vision-language RL training compared to relying solely on answer rewards. In standard RL training, the objective is driven only by the correctness of the final answer $a$:
$\nabla_{\theta}\,\mathbb{E}_{s\sim\pi_{\theta}}[r_{\text{ans}}(a,a^{*})],$ (5)
where $s=(t,a)$ is the trajectory with reasoning trace $t$ and final answer $a$. Since $r_{\text{ans}}$ depends solely on the match between $a$ and the ground truth $a^{*}$, the intermediate reasoning process $t$ receives no direct supervision. For VLMs, this is particularly problematic: the LLM backbone is far stronger than the vision encoder, and thus dominates the generation of $t$. With continued RL training under a single reward, gradients become high-variance and the model learns to exploit language priors – achieving correct answers without grounding in the visual input.
Reward decomposition.
We instead optimize a joint objective
$J(\theta)\;=\;\mathbb{E}_{s\sim\pi_{\theta}}\!\left[\,r_{\text{visual}}(c,x)\;+\;\,r_{\text{ans}}(a,a^{*})\right],$ (6)
where $r_{\text{visual}}$ measures informativeness of the visual perception caption $c$ of visual input $x$, and $r_{\text{ans}}$ is the final answer reward. The resulting gradient
$\nabla_{\theta}J(\theta)\;=\;\alpha\,\nabla_{\theta}r_{\text{cap}}\;+\;\beta\,\nabla_{\theta}r_{\text{ans}}$ (7)
anchors updates to both visual perception and language reasoning modules, providing direct signals for both perception and language reasoning.
From an information-theoretic perspective, let $I$ denote the visual input, $Q$ the question, $C$ the visual perception representation, and $A$ the final answer. Mutual information $I(U;V)$ measures how much knowing $U$ reduces uncertainty about $V$ (Shannon, 1948). If training relies only on $r_{\text{ans}}$, the model primarily maximizes I(A;Q)I(A;Q)—making answers strongly dependent on the question—while neglecting $I(A;I)$, the dependence of answers on the visual input. This permits shortcut solutions that bypass perception. By additionally optimizing $r_{\text{cap}}$, which enforces high $I(C;I)$, the model strengthens the path from $I$ to $A$, thereby increasing $I(A;I)$ and ensuring that answers remain grounded in visual perception rather than language-only correlations.
2.4 Data Preparation
Vision-SR1-47K. Our RL dataset consists of approximately 47K examples collected from 24 open-source VLM benchmarks. It spans three key reasoning domains (Figure 1): mathematical reasoning (30.5%), which strengthens quantitative and logical abilities; commonsense knowledge (30%); and general visual understanding (39.5%), which grounds the model in visual question answering.
Table 1: Vision-SR1-47K data comprises three domains—Math, Knowledge, and General Visual Reasoning—providing diverse supervision for VLM generalization and adaptation.
Category Included Datasets Size (%) Math CLEVR-Math, GeoQA+, UniGeo, GEOS, Geometry3K, Super-CLEVR 14K 30.5% Science Knowledge TQA, ScienceQA, AI2D, PMC-VQA, VQA-RAD, EXAMS-V-train 14K 30% General Visual Reasoning ChartQA, DVQA, PlotQA, FigureQA, MapQA, TabMWP, A-OKVQA, IconQA, visual7w, OpenSpaces, Spacellava 18K 39.5%
Figure 2: We prompt Qwen-2.5-VL-7B to create the SFT cold-start dataset to learn the ideal format to get a quick start in the RL stage to reduce training steps and training time. Our cold start SFT dataset is the intersection of three source datasets, creating a final set of 9K examples. Our filtration process in Sec 2.4 guarantees zero false positives across text-only, caption-based, and direct CoT reasoning.
[…]
3 Experiments
To implement our Vision-SR1, we employ Qwen-2.5-VL-3B and 7B as base models. We use SFT to train our the base model on the SFT cold start dataset to learn the ideal format. Following the initial SFT stage, we train the model with GRPO. The RL phase is trained for 1 epoch on the Vision-SR1-47K dataset. During training, the policy model first generates visual perceptions from the input image, then produces language reasoning and final answer. Before computing the final GRPO advantage, we compute a self-reward for visual perception. This reward is derived from the policy model’s ability to answer the question using only its own generated visual perceptions, without relying on the input image $i$.
3.1 Baseline Methods
Vision-R1 (Huang et al., 2025): The first R1-style reinforcement learning approach, which relies solely on answer rewards as the training signal. However, since the original Vision-R1 was trained only on math-domain data and performs poorly on general-domain reasoning, we reproduce it using our 47K dataset to ensure a fair comparison.
Perception-R1 (Xiao et al., 2025): Similar in training style to Vision-R1, but incorporates pre-extracted visual annotations as an additional reward signal. These visual annotations are derived from a state-of-the-art proprietary multimodal LLM (not specified in the paper).
Visionary-R1 (Xia et al., 2025): Trained to produce a caption–reason–answer output format during RL, where the supervision signal comes from an external text-only LLM (not specified in the paper).
For fair comparisons, we only re-train Vision-R1 on our 47K dataset, since both Perception-R1 and Visionary-R1 require access to external annotations or supervision signals, which are undisclosed.
3.2 Benchmarks and Metrics
Our evaluation covers three areas to evaluate VLMs abilities. Specifically, the domains include (1) general visual understanding, (2) multimodal math reasoning (3) visual hallucination detection.
General Visual Understanding. We evaluate general visual understanding across five diverse benchmarks. MMMU (Yue et al., 2024) tests cross-modal reasoning and subject knowledge with 11.5K college-level, four-choice questions spanning six disciplines. MMMU-Pro (Yue et al., 2025) increases the difficulty with ten choices per question and adds a challenging vision-only setting, where all text is embedded within the image to necessitate robust visual parsing. MM-Vet (Yu et al., 2024b) assesses a broad range of integrated vision-language skills – including recognition, OCR, and math – using a unified LLM-based evaluation score. RealWorldQA (xAI, 2024) features $\sim$700 real-world images from vehicle captures, paired with spatially grounded questions that require verifiable answers. VisNumBench (Weng et al., 2025) specifically targets visual number sense through $\sim$1.9K questions covering seven numerical attributes and four estimation tasks.
Multimodal Mathematical Reasoning. We assess mathematical reasoning using two specialized benchmarks. MathVerse (Zhang et al., 2024a) consists of 2.6K diagram-centric problems (e.g., geometry, functions), each rendered in six visual-text variants to disentangle true visual understanding from linguistic shortcuts. Evaluation is based on step-by-step Chain-of-Thought (CoT) correctness. MATH-Vision (Wang et al., 2024) presents $\sim$3K competition-grade problems across 16 disciplines and five difficulty levels, stressing advanced multimodal reasoning.
Hallucination Diagnosis. To diagnose model failures, we use HallusionBench (Guan et al., 2024), a benchmark designed to pinpoint specific errors: (i) language-side hallucination, where visual context is ignored, and (ii) visual-illusion errors, where the image is misinterpreted. The benchmark’s binary yes/no format enables precise error analysis.
For our evaluations, we all use Gemini-2.5-flash (Comanici et al., 2025) to judge response correctness on non-multiple choice format question, serving as a proxy for human judgment.
3.3 Experimental Results
Table 2: Vision-SR1 vs. baselines. For Vision-R1, as noted in Section 3.1, the original model checkpoint was trained only on math-domain data. So we also reproduce it using our 47K dataset.
General Visual Understanding Visual Math & Hallucination
Methods MMMU MMMU $M$ RealWorld VisNum Math MATH Hallusion Avg.
-Pro -Vet QA Bench Verse -Vision Bench
Visionary-R1 (3B) by Xia et al. (2025) 27.4 30.6 63.8 56.9 10.0 45.0 36.5 30.0 37.5
Percention-R1 (7B) by Xiao et al. (2025) 36.8 40.9 78.4 69.4 15.9 57.6 41.2 65.4 50.7
Vision-R1 (7B) by Huang et al. (2025) 34.9 42.8 70.2 60.1 33.0 57.3 42.9 65.2 50.8
Backbone model: Qwen2.5-VL-3B
Zero-shot Inference (before RL) 30.5 25.5 68.4 65.4 15.7 44.3 35.8 40.9 40.8
Supervised Fine-tuning (before RL) 39.5 49.4 70.2 62.1 29.3 47.8 39.9 68.1 50.8
Vision-R1 47K data (fair comparison) 40.3 49.5 63.3 63.0 36.7 45.5 38.8 67.4 50.6
Vision-SR1 (our method) 40.8 49.6 69.7 66.1 41.9 48.5 38.5 68.3 52.9
Backbone model: Qwen2.5-VL-7B
Zero-shot Inference (before RL) 34.2 33.5 67.1 68.5 21.4 49.2 41.1 51.7 45.8
Supervised Fine-tuning (before RL) 41.8 51.8 79.4 65.6 35.7 55.5 42.8 67.8 55.1
Vision-R1 47K data (fair comparison) 47.7 54.8 78.9 69.9 39.4 54.7 46.0 67.6 57.4
Vision-SR1 (our method) 49.1 57.2 76.2 71.6 42.6 56.5 46.7 69.8 58.8
3.3.1 Vision-SR1 v.s. Baseline Methods
Table 2 presents a comprehensive comparison of Vision-SR1 with several baseline methods across diverse vision-language benchmarks. For example, with the Qwen2.5-VL-7B backbone, Vision-SR1 reaches 49.1 on MMMU-Pro and 57.2 on MMMU, outperforming Vision-R1 fair comparison runs (47.7 and 54.8, respectively). When averaged across all benchmarks, Vision-SR1 establishes a clear margin of improvement. With the 7B backbone, it achieves an average score of 58.8, compared to 57.4 for Vision-R1 and 55.1 for supervised fine-tuning. Even with the smaller 3B backbone, Vision-SR1 attains 52.9 average, outperforming all comparable baselines. These results demonstrate that Vision-SR1 significantly outperforms prior baseline models across both general-purpose and math-specific visual reasoning tasks, validating the effectiveness of our approach.
3.3.2 Ablation Study on Self-Reward
We train a control version of our model without the visual perception self-reward (Vision-SR1 w/o self-reward). This ablated model still follows a structured output (visual perception, CoT reasoning, and answer) but is optimized only with answer reward and format rewards. The self-visual reward for self-evaluating visual perception is removed. We note the only difference between Vision-SR1 w/o self-reward and Vision-R1 (Huang et al., 2025) lies in the output structure, i.e., using different system prompts, while all supervision signals (answer reward and format rewards) remain the same. Interestingly, our system prompt yields slightly better performance (+0.3 on average). Table 3 reports the ablation results. We find that removing the visual perception reward consistently leads to worse overall performance compared to including self-reward in the training process.
Table 3: Results of ablation study: Vision-SR1 v.s. Vision-SR1 without visual perception self-reward.
General Visual Understanding Visual Math & Hallucination
Methods MMMU MMMU $M$ RealWorld VisNum Math MATH Hallusion Avg.
-Pro -Vet QA Bench Verse -Vision Bench
Vision-SR1 (3B) 40.8 49.6 69.7 66.1 41.9 48.5 38.5 68.3 52.9
  $\vdash$ w/o self-reward 40.0 48.0 67.4 62.6 41.6 47.7 38.9 65.8 51.5
Vision-SR1 (7B) 49.1 57.2 76.2 71.6 42.6 56.5 46.7 69.8 58.8
  $\vdash$ w/o self-reward 48.8 55.3 78.4 70.9 41.4 54.8 45.3 66.4 57.7
3.3.3 Analysis on Language Shortcut
Table 4: Language Shortcut Rate (LSR) across different benchmarks. Lower values indicate better performance, as a reduced LSR reflects fewer language shortcuts during reasoning.
General Visual Understanding Visual Math & Hallucination
Methods MMMU MMMU $M$ RealWorld VisNum Math MATH Hallusion Avg.
-Pro -Vet QA Bench Verse -Vision Bench
Vision-SR1 (3B) 6.49 9.0 6.4 0.07 5.4 10.3 8.3 0.10 5.6
  $\vdash$ w/o self-reward 7.36 8.0 5.1 0.07 4.2 11.4 9.2 0.09 7.6
Vision-SR1 (7B) 9.12 8.0 8.3 0.07 10.9 4.7 13.2 0.07 6.7
  $\vdash$ w/o self-reward 8.86 7.0 7.3 0.08 10.2 15.5 13.4 0.09 7.9
We also introduce the Language Shortcut Rate (LSR), a metric designed to quantify how often a model produces the correct answer with an incorrect visual perception. A high LSR suggests the model is leveraging language knowledge prior rather than genuine visual understanding.
Our evaluation, follows a two-step process and uses Gemini-2.5-flash as a judge: (1) Visual Perception Extraction: for each model output, we extracted the generated visual perception, denoted as $\hat{C}$. (2) Self-Containment Check: we then provide the $\hat{C}$ and the original question $Q$ to Gemini-2.5-Flash evaluator. If the evaluator can reproduce the correct ground-truth answer using only this information, $\hat{C}$ is deemed self-contained. Based on this process, we define the metrics: The Language Shortcut Rate (LSR) is defined as the percentage of instances where the model produces an incorrect (not self-contained) visual perception but still gives the correct final answer:
$\text{LSR}=\frac{\#\{\text{incorrect visual perception{} \& correct answer}\}}{\#\{\text{total samples}\}}$
A higher LSR indicates that the model is answering correctly while bypassing visual perception, suggesting reliance on language prior shortcuts. An LSR of 0 indicates no shortcutting, i.e., every correct answer is supported by a correct, self-contained visual perception.
We compute the LSR for 7B model w/ and w/o self rewards on seven selected benchmarks for demo example in Table 4. An important finding is that the visual shortcut is the highest in multimodal mathematical reasoning, which raises important question to previous work R1-VL (Zhang et al., 2025), VLM-R1 (Shen et al., 2025), Vision-R1 (Huang et al., 2025): is multimodal RL training on mathematical datasets truly improving VLMs’ abilities to perform visual reasoning, or simply awakes the models’ language reasoning ability to guess without actually looking at visual information?
3.3.4 Analysis on Text-only Reasoning
An interesting question is how different training strategies affect the text-only reasoning capabilities of VLMs. In particular, by decoupling visual perception and language reasoning with two separate rewards, we ask whether these abilities can mutually reinforce one another. To examine this, we evaluated the text-only performance of VLMs after RL fine-tuning on multimodal data.
Specifically, we tested on four text-only datasets: MMLU-Pro and SuperGPQA (multi-disciplinary, general-domain benchmarks), and MATH-500 and GSM8K (mathematical reasoning tasks). Our results (Table 5) compare Vision-R1, our method, and pre-RL training baselines.
Table 5: Through self-reward, the model is implicitly rewarded for text-only reasoning, leading to improved performance in general reasoning and reduced degradation in math reasoning benchmarks.
Model MMLU-Pro SuperGPQA GSM8K MATH-500
Backbone model: Qwen2.5-VL-3B
Before RL 34.3 15.1 78.5 65.2
Vision-R1 47.7 23.1 82.2 66.0
Vision-SR1 48.1 23.2 83.2 68.6
Backbone model: Qwen2.5-VL-7B
Before RL 33.4 17.1 86.0 73.4
Vision-R1 53.4 26.7 85.5 68.2
Vision-SR1 56.1 26.3 87.6 70.8
First, we observe that on GSM8K and MATH-500, multimodal RL training, including both Vision-R1 and our method, degrades text-only reasoning performance. This observation aligns with recent findings on “text-only forgetting” in VLMs Zhang et al. (2024b); Ratzlaff et al. (2025), which show that visual instruction tuning can impair language reasoning (particularly in mathematics) depending on the underlying LLM. Second, compared to Vision-R1, our method proved more effective at mitigating performance degradation on text-only mathematical benchmarks (MATH-500, GMS8K) and yielded larger gains on general knowledge tasks (MMLU-Pro, SuperGPQA). This indicates that separating the optimization signals for visual perception and language reasoning helps preserve text-only competencies, while still enabling improvements from multimodal training.
4 Related Work
4.1 Post-Training Vision-Language Models
Recent vision-language models have increasingly leveraged post-training alignment techniques, including instruction tuning and reinforcement learning, to enhance general-purpose multimodal performance (Liu et al., 2023b; Bai et al., 2025; Chen et al., 2024; et al, 2024; Huang et al., 2025). For example, LLaVA (Liu et al., 2023b) is tuned on GPT-4 generated (image, question, answer) pairs, coupling a CLIP encoder with Vicuna to produce a visual chat assistant that imitates some GPT-4 vision capabilities. InstructBLIP (Dai et al., 2023) introduces an instruction-aware query transformer tuned on 26 datasets, which yields a model that substantially outperforms even larger models on zero-shot benchmarks. Beyond standard instruction-tuning methods like LLaVA and InstructBLIP, recent work increasingly uses reinforcement learning (RL) to align vision-language models for better reasoning (Huang et al., 2025; Xia et al., 2025; Xiao et al., 2025). Many of these methods, inspired by techniques from DeepSeek-R1 (DeepSeek-AI et al., 2025), focus on sophisticated reward engineering. Strategies include providing step-wise rewards to supervise the intermediate reasoning (Zhang et al., 2025), adding explicit visual annotations to ground truth for calculating visual rewards (Xiao et al., 2025), and applying RL in a two-stage curriculum that first strengthens text-only reasoning (Peng et al., 2025b). As a complementary approach, RL from AI Feedback for VLMs demonstrates that preference-based alignment is also a powerful signal, showing it can substantially reduce object hallucination by learning from AI-generated feedback (Yu et al., 2024a).
4.2 Self-Rewarding Reinforcement Learning
The existing reinforcement learning with verifiable rewards (RLVR) methods heavily rely on high-quality reward models or human feedback, creating a major bottleneck for scalability (Peng et al., 2025a; Dai et al., 2025; Li et al., 2025c; Luu et al., 2025). To overcome this, recent work explores self-rewarding approaches, where the model itself provides intrinsic reward signals during RL post-training, an idea first pioneered by Yuan et al. (2025). Building on self-rewarding language models, methods replace external reward models with the model’s own confidence and uncertainty (logit-based self-certainty) or self-verification of its solutions, and even elicit a latent endogenous reward already present inside base LLMs (Zhao et al., 2025; Li et al., 2025a; Simonds et al., 2025; Zheng et al., 2025; van Niekerk et al., 2025). For example, RLIF leverages self-certainty as a reward, achieving comparable performance to GRPO while improving out-of-distribution generalization (Zhao et al., 2025). Similarly, RLSC optimizes a self-confidence reward to secure large accuracy gains with only a few training samples (Li et al., 2025a).
Although self-generated reward signals have thrived in text-only LLMs, only a few works extend this idea to VLMs (Zhou et al., 2024; Lee et al., 2025; Holmes & Chi, 2025), largely due to the complexity of the visual modality and the difficulty of properly defining and evaluating reward signals that capture visual perception. Recent progress includes Calibrated Self-Rewarding, which iteratively generates candidates, self-scores them with step-wise, visually constrained rewards, and fine-tunes via direct preference optimization (DPO) (Zhou et al., 2024). Similarly, RG-VLM uses a VLM to directly label rewards for offline trajectories in long-horizon visual tasks, serving as an auxiliary signal that boosts generalization (Lee et al., 2025). Beyond judgment-based signals, ARES derives dense shaped rewards from attention weights to accelerate learning under sparse or delayed feedback (Holmes & Chi, 2025). These works show that internal visual signals can provide rich reward feedback for VLM alignment without costly supervision, yet the reward is not integrated end-to-end, where the policy receives both visual perception and answer rewards during training.
5 Conclusion and Future Work
In this paper, we introduce Vision-SR1, a self-rewarded reinforcement learning framework that decomposes vision-language understanding into visual perception and language reasoning by explicitly rewarding visual perception by VLM itself. Vision-SR1 strengthens visual perception and reduces language shortcuts, thereby improving VLM performance across several domains of vision-language tasks. Our proposed metric LSR further shows how perception reward lowers the tendency of models to answer via language shortcut rather than genuine visual reasoning.
This work opens up several future research directions:
First, future research can explore more explicit perception rewards such a rewarding visual embeddings in answering questions directly, rather than converting them into textual cues to reduce information loss and lead to more robust and grounded visual understanding.
Second, exploring self-evolving VLMs that reward both visual perception and answers without using external signals (e.g. human labels) that can enhance VLM’s general visual reasoning abilities.
It is also important to recognize that some of the observed mathematical gains from RL training in VLMs may come from spurious effects – for instance, recalibrating the LLM backbone’s output distribution can boost multimodal math performance without true visual grounding. This suggests that improvements in accuracy may sometimes reflect better exploitation of language shortcuts rather than genuine perception gains. Therefore, future work can also explore more analysis to disentangle visual grounding from shortcut learning. Establishing a stronger foundation for measuring and mitigating language shortcuts (benchmarks) will enable the community to design more principled training objectives and ultimately build VLMs with deeper, more reliable visual reasoning.
Acknowledgements
We thank Hongming Zhang, Linfeng Song, Ruosen Li, Zhiyuan Ren for their helpful discussions!
Appendix A Experiment Details
A.1 Prompt Templates
This section presents the prompt templates used for constructing the cold start training data and Model Training prompt. The See-Think prompt is used for generating SFT See-Think data and model training. The Caption-Reasoner prompt is used to generate text-only caption reasoner SFT data and self-reward during training.
[…]
A.2 LLM-as-a-Judge Prompt
We use Gemini-2.5-flash as our LLM-as-a-Judge to evaluate
LLM-as-a-Judge Prompt Template Model: Gemini-2.5-flash Prompt Message:
Question: {Question}
Reference: {Reference} Candidate: {Candidate} You are provided a question, a gold answer, and a candidate answer. Your task is to judge the correctness of the candidate answer. Return your judgment enclosed with <judgment></judgment><judgment></judgment>. Note: {Question} is a placeholder for the actual question; {Reference} is a placeholder for the gold answer; {Candidate} is a placeholder for the model response.